### v1

@@ 1 Introduction
In this paper, we use the Subsystem Dependency Recovery Approach (SyDRA), described in the work of Ullmann et al. (2023), to help game engine developers understand game engine architecture and therefore make informed game engine development choices. By applying SyDRA to 10 open-source game engines, we obtain architectural models that can be used to compare game engine architectures and identify and solve problems such as high coupling and low cohesion. Through a controlled experiment with 16 participants, we show that the inspection of the architectural models derived from SyDRA enables developers to complete tasks related to architectural understanding and impact analysis in less time and with higher correctness than without these models.
@@
  • RQ1: Do architectural models help developers understand game engines more effectively than by simply inspecting the code?
  • RQ2: Do architectural models help game engine developers perform game engine impact analysis more effectively than by simply inspecting the code?
@@ 2 Related Work
Agrahari and Chimalakonda (2021, p. 1) developed and used AC2, a software analysis tool [^8] https://github.com/dheerajVagavolu/AC2 “to generate call graphs and collaboration graphs across three releases” of Unreal Engine. These graphs helped them identify architectural patterns in components, observe their evolution and “aid in better comprehension of this complex and widely used game engine for researchers and practitioners”.
@@ 3 Approach
Our implementation of SyDRA is available on Zenodo [^9] https://zenodo.org/records/7970779, along with the results from our analysis of 10 open-source game engines, which we also show and discuss. We obtain these results by implementing SyDRA’s steps as follows:
@@
FES Front-End Manages GUI, menus, heads-up display (HUD), and video playback.
@@
5. Architectural Model Generation: For each selected game engine, we used the data obtained from Steps 3 and 4 to generate an architectural model. These models were then loaded into Moose, a software analysis platform which we describe in more detail in Section 5.3. This step is done semi-automatically.
6. Architectural Model Visualisation: For each selected game engine, we use the “Architectural map” visualisation from Moose to generate a visual representation of the include graph, files, folders and subsystems. This step is done semi-automatically.
@@
Additionally, based on the number of nodes and edges on each include graph, we computed the following metrics, illustrated in Figure 2:
@@
In Section 4, we describe the structure of the architectural models obtained by applying SyDRA to 10 open-source game engines. By observing the subsystems and relationships among subsystems, we can explain the game engines they represent, and how game engine developers can use this information to increase cohesion through folder reorganization and decreasing coupling. Furthermore, in Section 5, we show examples of how the use of architectural models can inform developers performing architecture understanding and impact analysis tasks.
@@ 4 Application
As a result of applying the SyDRA, we obtained 10 architectural models, one for each of the selected game engines. These models are graphs where each node is a file, and each edge is an include relationship between files. For example, using Moose “Architectural map”, these files can be viewed clustered by subsystem, or individually, as shown in Figure 3.
@@
In this section, we show some examples of visual and numerical information extracted from these 10 architectural models and how they help developers understand three aspects of game engines: subsystem coupling, subsystem cohesion and coupling between files and subsystems.
@@ 4.1 Subsystem Coupling
The models produced by the SyDRA approach enable us to visualize the high-level architecture of the game engine, meaning it shows us what features are available and how they depend on each other. In Unreal Engine and Godot, we detected all 16 subsystems described in the reference architecture (see Figure 4). In the remaining game engines, we detected 12 or more subsystems. The only exception was OlcPixelGameEngine, in which we detected only five subsystems. Therefore, 90% of the game engines we analysed contain at least 75% of the subsystems described in the reference architecture, which shows there are many similarities between the reference architecture and the actual architecture of open-source game engines.
@@
Pair Count Pair Count
@@ 4.2 Subsystem Cohesion
We demonstrate this folder organisation pattern in Figure 6. While its rationale is not explained in O3DE’s documentation, we believe it was created to break down the Platform Independence Layer (PLA), a large subsystem that encompasses many features, into folders with a lower file count. However, while highlighting the separation between O3DE’s core and other subsystems, this organisation does not use its folder hierarchy to cohesively group all subsystem code under one folder. For example,AzQtComponents, which is a part of the Front-End (FES) subsystem, is in the same hierarchy level as AzFramework, a folder which contains files for several subsystems.
@@
While at first glance such folder-level refactoring may appear merely aesthetic, it can help developers re-purpose game engines. By eliminating unnecessary nesting and clustering files by subsystem, developers can focus on going forward with the development new features without being encumbered by legacy architectural structures and naming that no longer serve a purpose.
@@ 4.3 File and Subsystem Coupling
The inspection of the include relationships between files and the cross-referencing of this information with game engine documentation, when it exists, may give us detailed insights into how a particular part of a subsystem works.
@@
For example, if we look into Panda3D’s Low-Level Renderer (LLR) subsystem, we observe it has a high in-degree because of its display folder, which contains files implementing graphics-related functionality used by the Scene Graph / Culling Optimizations (SGC) and Visual Effects (VFX) subsystems.
One of these files, graphicsStateGuardian.h, implements a graphics state guardian (GSG), which receives high-level rendering instructions (e.g., drawing a character present in the scene graph) and then handles low-level rendering instructions in a format the operating system and graphics hardware can understand. As explained by Goslin and Mine (2004, p. 112):
@@ 5 Controlled Experiment
Variable X Variable Y - Tool
Run VS Code Moose + VS Code
@@ 5.3 Materials
In this study, participants analysed Godot [^14] https://github.com/godotengine/godot, a cross-platform, free and open-source game engine released by Juan Linietsky and Ariel Manzur in 2014. We chose Godot due to its relevance to the open-source developer community on GitHub, as explained in Section 3. While control group participants used exclusively Visual Studio Code to analyse Godot’s source code, treatment group participants used both Moose + Visual Studio Code, and therefore had access to the “Architectural map” visualisation of Godot produced with SyDRA.
We created instructional documents to teach participants to use the given tools. These documents provide step-by-step instructions accompanied by screenshots, illustrating how to use tool features. Additionally, at the end of each document, there were three optional exercises to motivate participants to put into practice the instructions they just read. As we explain in Section 5.5, we asked participants to read the document related to the tool they were about to use before performing the tasks.
@@
In this section, we provide an overview of the main features of Moose and Visual Studio Code and our rationale for choosing these tools for the study. We also describe the features participants used during the experiment to aid them in the completion of the tasks and show examples of the screenshots included in the instructional documents. Moose: A platform for software analysis [^15] https://moosetechnology.org/ composed of several tools built on top of the Pharo [^16] https://pharo.org/ programming language. It enables users to define metamodels and create models based on them. It also allows users to visualise these models as “Architectural Maps”, as explained in Section 3. Model entities can be written (or “propagated” in Moose’s jargon) to a bus, which is a channel of communication between tools (Anquetil et al., 2020, p.130). Moose tools, such as the “Architectural Map”, can then read entities from the bus and do something with them (e.g., draw a visualisation).
Figure 10: Moose with “Architectural Map” visible on the top right, used by the treatment group participants.
@@
During the study, as shown in Figure 10, treatment group participants located files and folders using Moose “Architectural Map” (top right) and propagated them to a built-in source code browser to inspect the source code (bottom). Participants could launch Visual Studio Code from Moose built-in editor to use features such as code folding, syntax highlighting, and search, not available on Moose. They could also use Moose tag browser to see the list of files clustered into each subsystem, represented as coloured tags (top left). Visual Studio Code: A source code editor [^17] https://code.visualstudio.com/ released by Microsoft in 2015. Also commonly referred to as VS Code. During the study, control group participants used it to locate files and folders inside the Godot repository, read source code and search for words as directed by the task statements, as we show in Figure 11. We chose VS Code due to its broad popularity among software developers. According to the Stack Overflow Developer Survey 2023, 74.09% of professional developers and 78.39% of developers learning to code use VS Code [^18] https://survey.stackoverflow.co/2023/#technology-most-popular-technologies.
@@ 5.4 Tasks
Suppose the rich text functionality in the Front-end subsystem of Godot was removed. Please mention all files which may have to be changed as a result of the removal of these functionalities. Suppose the rich text functionality in the Front end subsystem of Godot was removed. Please mention all files which may have to be changed as a result of the removal of these functionalities.
We wrote our task statements based on those provided in Appendix A of the original experiment (Briand et al., 2001, p. 527). We changed the statements slightly to make them easier for novice developers to understand. We also adapted task statements to reflect the steps participants in different groups had to perform to find files in the tools they were using, as demonstrated in Table 4. For example, while in Task 2, we asked treatment group participants to “Expand the Audio subsystem” and then “propagate” a given file, we asked control group participants to search the file by name and then open it [^19] See Folder 2 of the replication package for the complete task statements and the list of modifications we made to the original task statements.. This way, we ensured both participants were directed to the same file, even though they followed different steps to find and open it.
@@ 5.5 Procedures
The debriefing questionnaire was divided into two parts [^20] See Folder 2 of the replication package for the complete debriefing questions and Folder 4 for the participant’s answers.. In questions 1 to 7, we asked participants about their professional backgrounds and demographics. In questions 8 to 13, we asked participants to make a workload assessment of the tasks they performed based on the NASA TLX (Task Load Index) questionnaire. We chose NASA TLX because it has been used for over 20 years by several studies that evaluate software development (Al Madi et al., 2022, p.668), interface design and decision-making activities (Hart, 2006, p.906).
@@ 5.7 Data Analysis Procedures
We collected data from all participants during a single experimental run, which lasted about one month. Therefore, eight data points were available for the control group and eight for the treatment group. All participants answered all tasks and debriefing questions. In this section, we summarise the statistical techniques [^21] See Folder 3 of the replication package for statistical testing results. we used to determine whether the data collected during the study was statistically significant:
@@
    Effect Size: We computed the effect size for each variable by calculating the difference between the control group and treatment group arithmetic means, then dividing the result by the geometric mean of the control group and treatment group standard deviations [^23] https://meera.seas.umich.edu/power-analysis-statistical-significance-effect-size.html.
@@ 5.8 Results
In impact analysis tasks, the control group completed tasks faster but also less correctly than the treatment group, as we show in Figure 12(b) and Figure 12(d). This happened because, by using exclusively VS Code, participants had more difficulty finding all outgoing and incoming include relationships between files and ended their analysis prematurely. In contrast, participants using SyDRA could succinctly visualize relationships as arrows in the “Architectural Map”, which helped them complete impact analysis more correctly.
@@
Variable $\overline{X}$ Control $\overline{X}$ Treatment Effect Size Z
@@ 7 Conclusion
Through a controlled experiment with 16 software developers, we evaluated the extent to which SyDRA helps developers perform architecture understanding and impact analysis tasks. We asked developers to analyse Godot, an open-source game engine. While control group participants used exclusively VS Code to analyse Godot’s source code, treatment group participants used both Moose + Visual Studio Code, and therefore had access to the “Architectural map” visualisation of Godot produced with SyDRA. Our experiment’s most important contribution is showing a small yet statistically significant decrease in completion time for architectural understanding tasks when using VS Code in conjunction with the SyDRA-generated architectural model.
Finally, our analysis of the answers in the experiment’s debriefing questionnaires shows that by using SyDRA developers can better understand game engine architecture while not increasing their perceived workload. In future work, we intend to ask game engine developers to evaluate our suggestions for subsystem coupling reduction (Section 4.1) and subsystem cohesion increase (as shown in Section 4.2), as well as the emergent game engine architecture we propose in Section 4.4. We aim to understand whether these suggestions are useful in a professional video game development setting. Moreover, we are aware SyDRA is flexible enough to be used in the analysis of other software families which are fundamental to video game production, such as image and sound editors and 3D modelling software. Therefore, we intend to apply SyDRA to these software families, to identify common architectural structures that can guide their creation, maintenance and evolution.

### v2

@@ 1 Introduction
In this paper, we use the Subsystem Dependency Recovery Approach (SyDRA), described in our previous work (Ullmann et al., 2023), to help game engine developers understand game engine architecture and therefore make informed game engine development choices. By applying SyDRA to 10 open-source game engines, we obtain architectural models that can be used to compare game engine architectures and identify and solve problems such as high coupling and low cohesion. Through a controlled experiment with 16 participants, we show that the inspection of the architectural models derived from SyDRA enables developers to complete tasks related to architectural understanding and impact analysis in less time and with higher correctness than without these models.
@@
  • RQ1: Does SyDRA help developers understand game engines more effectively than by simply inspecting the code?
  • RQ2: Does SyDRA help game engine developers perform game engine impact analysis more effectively than by simply inspecting the code?
@@ 2 Related Work
Agrahari and Chimalakonda (2021, p. 1) developed and used AC2, a software analysis tool “to generate call graphs and collaboration graphs across three releases” of Unreal Engine. These graphs helped them identify architectural patterns in components, observe their evolution and “aid in better comprehension of this complex and widely used game engine for researchers and practitioners”.
@@ 3 Approach
Our implementation of SyDRA is available on GitHub [^8] https://github.com/gamedev-studies/game-engine-analyser, along with data and visualisations resulting from our analysis of 10 open-source game engines, which we also show and discuss in previous work (Ullmann et al., 2023). We obtain these results by implementing SyDRA’s steps as follows:
@@
FES Front End Manages GUI, menus, heads-up display (HUD), and video playback.
@@
5. Architectural Model Generation: For each selected game engine, we use the data obtained from Steps 3 and 4 to generate an architectural model. We then load these models into Moose, a software analysis platform which we describe in more detail in Section 5.3. This step is done semi-automatically.
6. Architectural Model Visualisation: For each selected game engine, we use the visualisation titled “Architectural map” from Moose to generate a visual representation of the include graph, files, folders and subsystems. This step is done semi-automatically.
@@
Additionally, based on the number of nodes and edges on each include graph, we compute the following metrics, illustrated in Figure 2:
@@
It took us approximately six months to apply SyDRA’s steps to 10 open-source game engines. The most time-consuming step was subsystem detection (Step 3), performed exclusively by the first author over approximately two months. In Section 4, we present the architectural models derived from the application of SyDRA. Furthermore, in Section 5, we show examples of how the use of architectural models can inform developers performing architecture understanding and impact analysis tasks.
@@ 4 Application
As a result of applying the SyDRA, we obtained 10 architectural models, one for each of the selected game engines. These models are graphs where each node is a file, and each edge is an include relationship between files. For example, using the Architectural Map from Moose, these files can be viewed clustered by subsystem, or individually, as shown in Figure 3.
@@
Examining subsystems and their relationships in the Architectural Map serves as a starting point for game engine developers seeking to understand the structure and functionality of a game engine. In this section, we show some examples of visual and numerical information extracted from these 10 architectural models and how they help developers understand three aspects of game engines: subsystem coupling, subsystem cohesion and coupling between files and subsystems.
@@ 4.1 Subsystem Coupling
The models produced by the SyDRA enable us to visualize the high-level architecture of the game engine, meaning it shows us what features are available and how they depend on each other. In Unreal Engine and Godot, we detected all 16 subsystems described in the reference architecture (see Figure 4). In the remaining game engines, we detected 12 or more subsystems. The only exception was OlcPixelGameEngine, in which we detected only five subsystems. Therefore, 90% of the game engines we analysed contain at least 75% of the subsystems described in the reference architecture, which shows there are many similarities between the reference architecture and the actual architecture of open-source game engines.
@@
Pair Count Pair aCount
@@ 4.2 Subsystem Cohesion
We demonstrate this folder organisation pattern in Figure 6. While its rationale is not explained in O3DE’s documentation, we believe it was created to break down the Platform Independence Layer (PLA), a large subsystem that encompasses many features, into folders with a lower file count. However, while highlighting the separation between O3DE’s core and other subsystems, this organisation does not use its folder hierarchy to cohesively group all subsystem code under one folder. For example,AzQtComponents, which is a part of the Front End (FES) subsystem, is in the same hierarchy level as AzFramework, a folder which contains files for several subsystems.
@@
While at first glance such folder-level refactoring may appear merely aesthetic, it can help developers re-purpose game engines. By eliminating unnecessary nesting and clustering files by subsystem, developers can focus on going forward with the development of new features without being encumbered by legacy architectural structures and naming that no longer serve a purpose.
@@ 4.3 File and Subsystem Coupling
The inspection of the include relationships between files and the cross-referencing of this information with game engine documentation, when it exists, may give us detailed insights into how a particular part of a subsystem works. In this section, we inspect subsystems in Panda3D and GamePlay3d, as shown in Figure 8.
@@
For example, if we look into Panda3D’s Low-Level Renderer (LLR) subsystem, we observe it has a high in-degree because of its display folder, which contains files implementing graphics-related functionality used by the Scene Graph / Culling Optimizations (SGC) and Visual Effects (VFX) subsystems. For instance, one of these files, graphicsStateGuardian.h, implements a graphics state guardian (GSG), which receives high-level rendering instructions (e.g., drawing a character present in the scene graph) and then handles low-level rendering instructions in a format the operating system and graphics hardware can understand. As explained by Goslin and Mine (2004, p. 112):
@@ 5 Controlled Experiment
aVariable Xa Variable Y - Tool
Run aVS Codea aMoose + VS Codea
@@ 5.3 Materials
In this study, participants analysed Godot [^13] https://github.com/godotengine/godot, a cross-platform, free and open-source game engine released by Juan Linietsky and Ariel Manzur in 2014. We chose Godot due to its relevance to the open-source developer community on GitHub, as explained in Section 3. While control group participants used exclusively Visual Studio Code to analyse Godot’s source code, treatment group participants used both Moose + Visual Studio Code, and therefore had access to the Architectural Map visualisation of Godot produced with SyDRA.
We created instructional documents to teach participants to use the given tools. As we explain in Section 5.5, we asked participants to read the document related to the tool they were about to use before performing the tasks. These documents provided an overview of features such as file searching and browsing, illustrated by screenshots. At the end of each document, we also provided three optional warm-up exercises designed to reinforce the instructions. Although similar to the tasks in the experiment, these warm-ups were simpler and focused on a single feature at a time. For example, in the Visual Studio Code warm-ups, we asked participants to use the search feature to find and count the number of files containing the word “Music”. They could then verify their answers on the last page of the document.
@@
In this section, we provide an overview of the main features of Moose and Visual Studio Code and our rationale for choosing these tools for the study. We also describe the features participants used during the experiment to aid them in the completion of the tasks and show examples of the screenshots included in the instructional documents.
Moose: A platform for software analysis [^14] https://moosetechnology.org/ composed of several tools built on top of the Pharo [^15] https://pharo.org/ programming language. It enables users to define metamodels and create models based on them. It also allows users to visualise these models as Architectural Maps, as explained in Section 3. Model entities can be written (or “propagated” in Moose’s jargon) to a bus, which is a channel of communication between tools (Anquetil et al., 2020, p.130). Moose tools, such as the Architectural Map, can then read entities from the bus and do something with them (e.g., draw a visualisation).
Figure 10: Moose with Architectural Map visible on the top right.
@@
During the study, as shown in Figure 10, treatment group participants located files and folders using the Architectural Map (top right) and propagated them to a built-in source code browser to inspect the source code (bottom). Participants could launch Visual Studio Code from Moose built-in editor to use features such as code folding, syntax highlighting, and search, not available on Moose. They could also use Moose tag browser to see the list of files clustered into each subsystem, represented as coloured tags (top left).
Visual Studio Code: A source code editor [^16] https://code.visualstudio.com/ released by Microsoft in 2015. Also commonly referred to as VS Code. During the study, control group participants used it to locate files and folders inside the Godot repository, read source code and search for words as directed by the task statements, as we show in Figure 11. We chose VS Code due to its broad popularity among software developers. According to the Stack Overflow Developer Survey 2023, 74.09% of professional developers and 78.39% of developers learning to code use VS Code [^17] https://survey.stackoverflow.co/2023/#technology-most-popular-technologies.
@@ 5.4 Tasks
Suppose the rich text functionality in the Front end subsystem of Godot was removed. Please mention all files which may have to be changed as a result of the removal of these functionalities. Suppose the rich text functionality in the Front end subsystem of Godot was removed. Please mention all files which may have to be changed as a result of the removal of these functionalities.
We wrote our task statements based on those provided in Appendix A of the original experiment (Briand et al., 2001, p. 527). We changed the statements slightly to make them easier for novice developers to understand. We also adapted task statements to reflect the steps participants in different groups had to perform to find files in the tools they were using, as demonstrated in Table 4. For example, while in Task 2, we asked treatment group participants to “Expand the Audio subsystem” and then “propagate” a given file, we asked control group participants to search the file by name and then open it [^18] See Folder 2 of the replication package for the complete task statements.. This way, we ensured both participants were directed to the same file, even though they followed different steps to find and open it.
@@ 5.5 Procedures
The debriefing questionnaire was divided into two parts [^19] See Folder 2 of the replication package for the complete debriefing questions/answers.. In questions 1 to 7, we asked participants about their professional backgrounds and demographics. In questions 8 to 13, we asked participants to make a workload assessment of the tasks they performed based on the NASA TLX (Task Load Index) questionnaire. We chose NASA TLX because it has been used for over 20 years by several studies that evaluate software development (Al Madi et al., 2022, p.668), interface design and decision-making activities (Hart, 2006, p.906).
@@ 5.7 Data Analysis Procedures
We collected data from participants throughout 16 sessions spread across one month. Therefore, eight data points were available for the control group and eight for the treatment group. All participants answered all tasks and debriefing questions. We used the following statistical techniques [^20] See Folder 3 of the replication package for statistical testing results. to determine whether the data collected during the study was statistically significant:
@@
    Effect Size: We computed the effect size for each variable by calculating the difference between the control group and treatment group arithmetic means, then dividing the result by the geometric mean of the control group and treatment group standard deviations.
@@ 5.8 Results
In impact analysis tasks, the control group completed tasks faster but also less correctly than the treatment group, as we show in Figure 12(b) and Figure 12(d). This happened because, by using exclusively VS Code, participants had more difficulty finding all outgoing and incoming include relationships between files and ended their analysis prematurely. In contrast, participants using SyDRA could succinctly visualize relationships as arrows in the Architectural Map, which helped them complete impact analysis more correctly.
@@
[…]
@@ 7 Conclusion
Through a controlled experiment with 16 software developers, we evaluated the extent to which SyDRA helps developers perform architecture understanding and impact analysis tasks. We asked developers to analyse Godot, an open-source game engine. While control group participants used exclusively VS Code to analyse Godot’s source code, treatment group participants used both Moose + Visual Studio Code, and therefore had access to the Architectural Map visualisation of Godot produced with SyDRA. Our experiment’s most important contribution is showing a small yet statistically significant decrease in completion time for architectural understanding tasks when using VS Code in conjunction with the SyDRA-generated architectural model.
Finally, our analysis of the answers in the experiment’s debriefing questionnaires shows that by using SyDRA developers can better understand game engine architecture while not increasing their perceived workload. In future work, we intend to ask game engine developers to evaluate our suggestions for subsystem coupling reduction (Section 4.1) and subsystem cohesion increase (as shown in Section 4.2), as well as the emergent game engine architecture we propose in Section 4.4. We aim to understand whether these suggestions are useful in a professional video game development setting, and we will seek collaboration with independent video game development companies to achieve this.
Moreover, we are aware SyDRA is flexible enough to be used in the analysis of other software families which are fundamental to video game production, such as image and sound editors and 3D modelling software. Therefore, we intend to apply SyDRA to these software families, to identify common architectural structures that can guide their creation, maintenance and evolution.